I See The Light Projected Balance Sheet As of December 31, 20x1 $ 34,710.00 67,500.00 Current Assets Cash Accounts ReceivableSee The Light Projected Income Statement For the Period Ending December 31, 20x1 $ 1.125,000.00 750,000.00 375,000.00 $ $ 23.The projected cost of a lamp is calculated based upon the projected increases or decreases to current costs. The present cost2 Materials Budget Lamp Kits Needed for Production Desired Ending Inventory Total Needed Less: Beginning Inventory Total PurcDivision N has decided to develop its budget based upon projected sales of 43,000 lamps at $46.00 per lamp. The company has r
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
I See The Light
Desired ending inventory of Finished Goods= 20% less than the opening quantity Note
Balance as of Dec 20X1        3,000.00 A
20% of balance           600.00 B=A*20%
Desired ending inventory        2,400.00 C=A-B
Answer 1 Answer No.
Production Budget Units Note
Projected Sales      43,000.00
Add: Desired ending inventory of Finished Goods        2,400.00 See C
Total Needed      45,400.00
Less: Beginning Inventory        3,000.00
Total Production     42,400.00 D Ans to 7.01
Answer 2
Materials budget
Lamp Kits Note
Total Production      42,400.00 See D
Number of Lamp Kits needed for producing 1 unit                1.00 E
Needed for production of lamps     42,400.00 F=D*E Ans to 8.01
Add: Desired ending inventory           750.00 Ans to 8.02
Total Needed     43,150.00 Ans to 8.03
Less: Beginning Inventory (opening balance)           500.00 Ans to 8.04
Total Purchases     42,650.00 G
Cost per piece             16.00 H Ans to 8.05
Cost of purchases 682,400.00 I=G*H Ans to 8.06
Calculation of Labor Costs Note
Current Labor Cost per lamp                2.00 J
Increase by 6%                0.12 K=J*6%
Revised Labor Cost per lamp                2.12 L=J+K
Production of lamps      42,400.00 See D Ans to 8.07
Total Labor Cost     89,888.00 M=L*D Ans to 8.08
Variable Factory Overhead Budget Note
Calculation of Variable Factory Overhead
Variable Factory Overhead per lamp                2.00 N
Increase by 4.5%                0.09 O=N*4.5%
Revised Variable Factory Overhead per lamp                2.09 P=N+O
Production of lamps      42,400.00 See D
Total Variable Factory Overhead     88,616.00 Q=P*D Ans to 8.09
Fixed Factory Overhead 285,000.00 R Ans to 8.10
Total Factory Overhead 373,616.00 S=Q+R Ans to 8.11